Two busses start at the same time towards each other from terminals A and B, 8 km apart. The time needed for the bus to travel A to B is 8 minutes, and of the second bus from B to A is 10 minutes. How much is the time needed by each bus to meet each if they travelled at their respective uniform speeds?
1
Expert's answer
2021-07-30T16:55:03-0400
"d=vt\\\frac{s}{t}=v_1+v_2=\frac{s}{t_1}+\frac{s}{t_2}\\
\frac{1}{t}=\frac{1}{8}+\frac{1}{10}\\t=4.44\ min"